OpenStudy (anonymous):

15+(1/3+3/5)*10 I need help solving this with step by step how to, I have the answer(out of book) but cant come up with the same

hero (hero):

Parentheses first...meaning perform the operations inside the parentheses first

hero (hero):

so add 1/3 + 3/5

hero (hero):

5/15 + 9/15 = 14/15

hero (hero):

Next multiply that by 10 to get 14/15 * 10 = 140/15

hero (hero):

now add 15 to that: 140/15 + 15

hero (hero):

I get 73/3 final answer

hero (hero):

What did the book say?

OpenStudy (anonymous):

140/15 + 15

hero (hero):

Oh

hero (hero):

Well, it can be simplified further to get 73/3 but 140/15 + 15 is good too

OpenStudy (anonymous):

the book says 24 1/3

OpenStudy (anonymous):

so I put 15/1 and multiply 140/15 and simplify?